OpenStudy (anonymous):

What is the 13th term in the Fibonacci Sequence? Term 1 2 3 4 5 6 7 8 9 Fibonacci Number 1 1 2 3 5 8 13 21 34

OpenStudy (king):

233?

OpenStudy (mathmate):

For larger values of n, you can also use the explicit formula: F(n)=(phi^n-psi^n)/sqrt(5) where phi=(1+sqrt(5))/2, and psi=(1-sqrt(5))/2 Here F(13)=233

OpenStudy (anonymous):

thank u for explaining

OpenStudy (anonymous):

can u show me how to fined my other question

OpenStudy (mathmate):

You're welcome! A note of caution: the formula involves irrational numbers, but it all cancels out in the end to give an integer. Using a calculator, it may involve rounding.